# EEG-Deformer: A Dense Convolutional Transformer for Brain-computer Interfaces

### v1

@@ IV-B Baselines
IV-B1 DGCNN
DGCNN employs a novel approach by utilizing a learnable adjacency matrix within its neural network training framework. This matrix dynamically captures and updates the relationships between different EEG channels throughout the training process. By adapting to the evolving data patterns, this method allows for a more flexible and accurate representation of the EEG signal.
IV-B2 LGGNet
LGGNet is a graph neural network inspired by neurology, designed to capture local-global-graph representations of EEG data. The input layer of LGGNet consists of temporal convolutions with multi-scale 1D convolutional kernels and attentive fusion, effectively capturing the temporal dynamics of EEG. This data feeds into local and global graph-filtering layers that model complex relationships within and between brain functional areas.
IV-B3 EEGNet
EEGNet is a streamlined convolutional neural network tailored for EEG-based BCIs. It incorporates depthwise and separable convolutions to develop an EEG-specific model that effectively integrates established EEG feature extraction techniques for BCI applications, making it a compact and effective EEG decoding neural network architecture.
IV-B4 TSception
TSception is a multi-scale convolutional neural network. The architecture of TSception integrates dynamic temporal layers, asymmetric spatial layers, and advanced fusion layers. This design synergistically extracts discriminative features by leveraging both temporal dynamics and spatial asymmetry across time and channel dimensions. This multifaceted approach enables TSception to effectively capture the complex patterns essential for accurate emotion recognition from EEG signals.
IV-B5 EEG-ViT
To explore the effects of combining CNNs and Transformers for decoding EEG signals, we adapt ViT [5], a purely Transformer-based architecture, to EEG data. We partition each EEG input into non-overlapping segments along the temporal dimension and then convert these segments into tokens using a linear projection layer. This allows ViT to effectively learn the temporal relationships among the tokens.
IV-B6 SSVEPformer
@@
IV-B7 EEG-Transformer
The EEG-Transformer is a convolutional Transformer that sequentially integrates CNN and Transformer architectures for EEG signal decoding. This model utilizes an EEGNet-based CNN feature extractor followed by a vanilla Transformer [23] to effectively process EEG signals.
IV-B8 EEG-Conformer
It has a similar design principle as EEG-Transformer, which combines CNN and Transformer to learn the spatial-temporal information encoded in EEG signals. EEG-Conformer utilizes one-dimensional CNNs to capture low-level local features and a self-attention module is directly connected to elucidate the global correlations inherent in these local temporal features.
TABLE III: Effects of different information purification methods on three benchmarking datasets, with the best results highlighted in bold.
Method Dataset I Dataset II Dataset III ACC std F1-macro std ACC std F1-macro std ACC std F1-macro std Using mean 77.35 6.99 76.99 7.33 73.82 12.13 71.72 12.93 68.82 15.44 65.47 18.90 Using std 77.54 7.54 77.04 8.07 73.94 14.42 72.34 14.77 72.27 15.20 69.19 19.00 Using power 82.72 8.00 82.36 8.52 79.32 7.87 75.83 11.54 73.18 15.63 69.99 19.80
TABLE IV: Effects of various information purification locations on three benchmarking datasets, with the best results highlighted in bold.
Method Dataset I Dataset II Dataset III ACC std F1-macro std ACC std F1-macro std ACC std F1-macro std At $\bm{F}_{i+1}$ 80.28 7.46 79.81 8.03 77.25 7.14 73.98 9.96 69.16 15.46 66.35 18.19 At $\bm{F}_{cg}^{i}$ 72.90 7.30 72.05 8.35 44.72 16.48 29.96 8.30 71.41 15.30 68.56 18.50 At $\bm{F}_{fg}^{i}$ 82.72 8.00 82.36 8.52 79.32 7.87 75.83 11.54 73.18 15.63 69.99 19.80
IV-C Experiment Settings
In this study, we implement generalized subject-independent settings, ensuring that test data information is not used during the training stage. We adopt a leave-one-subject-out (LOSO) approach for all three datasets. In each LOSO step, one subject’s data is set aside as test data. Of the remaining data, 80% is used for training and the remaining 20% serves as validation (development) data. We employ Accuracy (ACC) and Macro-F1 as our evaluation metrics which can be formalised as
@@ IV-C Experiment Settings
IV-D Implementation Details
The cross-entropy loss is utilized to guide the training. We use an Adam optimizer with an initial learning rate of 1e-3 and a weight decent of 1e-5. A cosine annealing schedule is adopted to adjust the learning rate during training. A dropout rate of 0.5 is applied on dataset I and II to avoid over-fitting. The dropout rate is changed to 0.25 on dataset III as it achieves a better performance on development set. The batch size is 64 for all the datasets.We train the model for 200 epochs, and the model with the best validation accuracy is used to evaluate the test data. The kernel lengths of CNN layers are calcuated by $0.1*f_{s}$, where $f_{s}$ is the sampling rate of the EEG segment. The odd length is need for Pytorch to achieve the same padding. As the $f_{s}$ of three datasets are 200 Hz, 128 Hz, and 500 Hz, the kernel lengths of CNN layers $l_{kernel}^{\textrm{I}}=21$, $l_{kernel}^{\textrm{II}}=13$, and $l_{kernel}^{\textrm{III}}=51$. We set the number of CNN kernels as 64 for all the datasets as [15]. We set $n_{head}=d_{atten}$, and tune it based on the performance on the validation (development) set. Hence, we have $n_{head}^{\textrm{I}}=32$ and $n_{head}^{\textrm{II}}=n_{head}^{\textrm{III}}=16$.
V Results and Analyses
V-A Classification Results
Attention Classification: The results for attention classification using dataset I, as presented in Table I, reveal that the EEG-Deformer outperforms all baseline methods. It achieves the highest accuracy and macro-F1 score in attention classification with 82.72% accuracy and a 82.36% macro-F1 score. Compared to EEG-Conformer, the next best performer, EEG-Deformer shows an improvement of 2.92% in accuracy and 3.35% in macro-F1 score. The results highlight that CNN-based methods generally outperform GNN-based ones and suggest that a combination of CNN and Transformer architectures yields better performance than using either CNN or Transformer alone.
Fatigue Classification: The same observation appears in the fatigue decoding task using dataset II, as detailed in Table I, the EEG-Deformer demonstrates superior performance, leading with an accuracy of 79.32% and a macro-F1 score of 75.83%. These results surpass those of the EEG-Conformer by 4.96% in accuracy and 4.18% in macro-F1 score. The same trend is shown here that the CNN-based methods are better than GNNs. And using Transformer without CNN layers yield low classification results.
V-B Ablation
To investigate the individual contributions of the FTL, Dense connections of IP-Units, and IP-Unit, we conduct an ablation study by selectively removing these layers and observing their impact on classification outcomes across three datasets. The findings, detailed in Table II, indicate that the omission of the FTL had the most detrimental effect on performance overall. Additionally, the removal of the IP-Unit results in the second-largest decrease in performance on datasets I and II. Eliminating dense connections also leads to reductions in both classification accuracy and macro-F1 scores across all datasets. These results collectively imply that each module within the EEG-Deformer plays a crucial role, working in unison to enhance its predictive capabilities.
V-C Ablation of IP-Units
@@
V-D Effect of Different Types of IP-Unit
Table III illustrates the impact of different information purification methods, which calculate mean, standard deviation (std), and power along the feature dimension of $F_{fg}^{i}$, on the performance across three datasets. The comparison focuses on their effects on accuracy and macro-F1 score. Notably, the ’power’ method consistently yields the highest accuracy and macro-F1 scores across all datasets, outperforming the other two methods. These findings underscore the superiority of using the ’power’ method for information purification in this context.
V-E Effect of Different Locations of The IP-Unit
The study assesses the impact of varying the placement of IP-Unit within each HCT block, focusing on three specific locations in Figure 4 : at the fused output, $\bm{F}_{i+1}$; at the coarse-grained representation, $\bm{F}_{cg}^{i}$; and at the fine-grained representation, $\bm{F}_{fg}^{i}$. The results of this evaluation are presented in Table IV. These findings reveal that applying IP-Unit to the output of the FTL leads to the highest decoding performance across all three datasets. This outcome highlights the effectiveness of this particular placement for IP-Unit within the HCT block structure.
V-F The Parameters Comparison of EEG-Deformer and Other Baseline Methods.
@@ V-E Effect of Different Locations of The IP-Unit
V-G Visualization
In this study, saliency maps [30] are utilized to visualize the areas deemed most informative by the neural networks. Figures 5 to 9 display saliency maps for five representative subjects and the average of all subjects across three datasets, corresponding to attention, fatigue, and mental workload classification tasks. For enhanced visualization, these saliency maps are normalized to values between 0 and 1. The human brain can be divided into frontal, temporal, parietal, and occipital functional areas. According to Figure 5, the areas most informative for attention classification are primarily located in the frontal (Fp1, F1, and AFF6) and parietal (CP5, P7, and P4) regions, which are known to be related to cognitive attention [31]. Figure 8 shows that for fatigue classification, the neural network primarily focuses on the frontal (F7, F3, and FCz), temporal (T3 and T5), and parietal (P3) areas, consistent with [32]. In the case of mental workload classification, we find consistent findings with those reported in [33]: the frontal (Fz and Fp2) and parietal (P4) areas are more informative for neural networks, as shown in Figure 9. These visualizations align with known brain activity patterns related to these cognitive processes and demonstrate the neural network’s ability to identify relevant brain regions for each task.
VI Conclusion
@@
Acknowledgment
This work was supported by the RIE2020 AME Programmatic Fund, Singapore (No. A20G8b0102) and was partially supported by the MoE AcRF Tier 1 Project (No. RT01/21).

### v2

@@ 4.2 Baselines
4.2.1 DGCNN
DGCNN introduces a learnable adjacency matrix that dynamically captures and updates relationships between EEG channels during training, allowing for a more flexible and accurate representation of the EEG signal.
4.2.2 LGGNet
LGGNet, a neurology-inspired graph neural network, captures local-global-graph representations of EEG data. Its input layer uses multi-scale temporal convolutions with attentive fusion to capture EEG dynamics, followed by graph-filtering layers that model complex relationships within and between brain regions.
4.2.3 EEGNet
EEGNet is a compact convolutional neural network designed for EEG-based BCIs, using depth-wise and separable convolutions to efficiently integrate EEG feature extraction techniques, making it highly effective for EEG decoding.
4.2.4 TSception
TSception is a multi-scale convolutional neural network that combines dynamic temporal layers, asymmetric spatial layers, and advanced fusion to capture complex patterns in EEG signals. This design effectively extracts discriminative features for accurate emotion recognition.
4.2.5 EEG-ViT
To explore the effects of combining CNNs and Transformers for decoding EEG signals, we adapt ViT [7], a purely Transformer-based architecture, for EEG data. We partition the EEG into non-overlapping segments along the temporal dimension and convert these segments into tokens using a linear projection layer, enabling ViT to process EEG effectively.
4.2.6 SSVEPformer
@@
4.2.7 EEG-Transformer
EEG-Transformer is a convolutional Transformer that combines an EEGNet-based CNN feature extractor with a vanilla Transformer [30] to decode EEG signals effectively.
4.2.8 EEG-Conformer
EEG-Conformer, like EEG-Transformer, combines CNNs and Transformers to capture spatial-temporal information in EEG signals. It uses 1-D CNNs for local feature extraction and a self-attention module to reveal global correlations in these temporal features.
4.3 Experiment Settings
In this study, we implement generalized subject-independent settings, ensuring that test data information is not used during the training stage. We adopt a leave-one-subject-out (LOSO) approach for all three datasets. In each LOSO step, one subject’s data is set aside as test data. Of the remaining data, 80% is used for training and the remaining 20% serves as validation (development) data. We employ Accuracy (ACC) and Macro-F1 as our evaluation metrics which can be formularized as
@@ 4.3 Experiment Settings
4.4 Implementation Details
The PyTorch implementation is available on this website [^1] https://github.com/yi-ding-cs/EEG-Deformer. The cross-entropy loss is utilized to guide the training. We use an Adam optimizer with an initial learning rate of 1e-3 and a weight decent of 1e-5. A cosine annealing schedule is adopted to adjust the learning rate during training. A dropout rate of 0.5 is applied on dataset I and II to avoid over-fitting. The dropout rate is changed to 0.25 on dataset III as it achieves a better performance on development set. The batch size is 64 for all the datasets.We train the model for 200 epochs, and the model with the best validation accuracy is used to evaluate the test data. The kernel lengths of CNN layers are calculated by $0.1*f_{s}$, where $f_{s}$ is the sampling rate of the EEG segment. The odd length is need for PyTorch to achieve the same padding. As the $f_{s}$ of three datasets are 200 Hz, 128 Hz, and 500 Hz, the kernel lengths of CNN layers $l_{kernel}^{\textrm{I}}=21$, $l_{kernel}^{\textrm{II}}=13$, and $l_{kernel}^{\textrm{III}}=51$. We set the number of CNN kernels as 64 for all the datasets as [18]. We set $n_{head}=d_{atten}$, and tune it based on the performance on the validation (development) set. Hence, we have $n_{head}^{\textrm{I}}=32$ and $n_{head}^{\textrm{II}}=n_{head}^{\textrm{III}}=16$.
5 Results and Analyses
5.1 Classification Results
Attention Classification: The results for attention classification using dataset I, as presented in Table 1, reveal that the EEG-Deformer outperforms all baseline methods. It achieves the highest accuracy and macro-F1 score in attention classification with 82.72% accuracy and a 82.36% macro-F1 score. Compared to EEG-Conformer, the next best performer, EEG-Deformer shows an improvement of 2.92% in accuracy and 3.35% in macro-F1 score. The results highlight that CNN-based methods generally outperform GNN-based ones and suggest that a combination of CNN and Transformer architectures yields better performance than using either CNN or Transformer alone.
Fatigue Classification: The same observation appears in the fatigue decoding task using dataset II, as detailed in Table 1, the EEG-Deformer demonstrates superior performance, leading with an accuracy of 79.32% and a macro-F1 score of 75.83%. These results surpass those of the EEG-Conformer by 4.96% in accuracy and 4.18% in macro-F1 score. The same trend is shown here that the CNN-based methods are better than GNNs. And using Transformer without CNN layers yield low classification results.
5.2 Ablation
To investigate the individual contributions of the FTL, Dense connections of IP-Units, and IP-Unit, we conduct an ablation study by selectively removing these layers and observing their impact on classification outcomes across three datasets. The findings, detailed in Table 2, indicate that the omission of the FTL had the most detrimental effect on performance overall. Additionally, the removal of the IP-Unit results in the second-largest decrease in performance on datasets I and II. Eliminating dense connections also leads to reductions in both classification accuracy and macro-F1 scores across all datasets. These results collectively imply that each module within the EEG-Deformer plays a crucial role, working in unison to enhance its predictive capabilities.
@@ 5.2 Ablation
5.3 Ablation of IP-Units
5.4 Effect of Different Types of IP-Unit
Table 3 illustrates the impact of different information purification methods, which are mean, standard deviation (std), and power on the performance across three datasets. For each layer’s output, $\textbf{{F}}_{fg}^{i}\in\mathbb{R}^{k\times 0.5*l_{t}^{i}}$, we will calculate the mean, std on the last dimension of $\textbf{{F}}_{fg}^{i}$, resulting the purified information, $\textbf{{I}}_{i}\in\mathbb{R}^{k}$. The comparison focuses on their effects on accuracy and macro-F1 score. Notably, the ’power’ method consistently yields the highest accuracy and macro-F1 scores across all datasets, outperforming the other two methods. These findings underscore the superiority of using the ’power’ method for information purification in this context.
5.5 Effect of Different Locations of The IP-Unit
The study assesses the impact of varying the placement of IP-Unit within each HCT block, focusing on three specific locations in Figure 4 : at the fused output, $\bm{F}_{i+1}$; at the coarse-grained representation, $\bm{F}_{cg}^{i}$; and at the fine-grained representation, $\bm{F}_{fg}^{i}$. The results of this evaluation are presented in Table 4. These findings reveal that applying IP-Unit to the output of the FTL leads to the highest decoding performance across all three datasets. This outcome highlights the effectiveness of this particular placement for IP-Unit within the HCT block structure.
5.6 The Parameters Comparison of EEG-Deformer and Other Baseline Methods.
5.7 Visualization
In this study, we use saliency maps [39] to visualize the regions that the neural networks identify as most informative. Figures 5 to 7 show saliency maps for five randomly selected subjects, as well as the averaged map across all subjects, for three datasets corresponding to attention, fatigue, and mental workload classification tasks. To improve clarity, these saliency maps are normalized to a range between 0 and 1. The brain is functionally divided into frontal, temporal, parietal, and occipital regions. According to Figure 5, the areas most informative for attention classification are primarily located in the frontal (Fp1, F1, and AFF6) and parietal (CP5, P7, and P4) regions, which are known to be related to cognitive attention [40]. Figure 6 shows that for fatigue classification, the neural network primarily focuses on the frontal (F7, F3, and FCz), temporal (T3 and T5), and parietal (P3) areas, consistent with [41]. In the case of mental workload classification, we find consistent findings with those reported in [42]: the frontal (Fz and Fp2) and parietal (P4) areas are more informative for neural networks, as shown in Figure 7. These visualizations align with known brain activity patterns related to these cognitive processes and demonstrate the neural network’s ability to identify relevant brain regions for each task. As shown in the figures, the saliency maps display generally consistent patterns for each task, indicating that our model consistently learns from task-related brain regions.
5.8 Limitations and future work
While the model demonstrates strong performance on the selected tasks, the evaluation does not cover all commonly used tasks in EEG-based classification. Expanding the range of tasks could provide a more comprehensive assessment of the model’s generalizability across different applications. In some datasets, the standard deviation of the model’s performance is higher compared to certain baseline methods. This suggests that, although the model performs well on average, it may exhibit greater variability across different subjects or trials, potentially affecting its reliability in certain cases. In the future, efforts should focus on improving the model’s robustness.
6 Conclusion